HUBBLE SPACE TELESCOPE DAILY REPORT #1917 PERIOD COVERED: 0000Z (UTC) 07/10/97 - 0000Z (UTC) 07/11/97 Daily Status Report as of 192/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7659 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make dark calibration images in order to update the acquisition bad pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 7276 (Synoptic Monitoring of Seasonal Phenomena on Mars) The WF/PC-2 was used to make seasonal ultraviolet imaging of the planet Mars. This season includes the formation of the north polar cap and hood system, a process which is not well documented; the start of the classic dust storm season on Mars, when large dust storms have often occurred; and an active cycle of small dust storms in Hellas. These observations will complete a period of continuous monitoring of Mars which commenced in Cycle 6. The observations were completed as planned, and no problems were reported. 1.3 Completed Five Sets of STIS/CCD 7601 (CCD Dark and Bias Monitor) The Space Telescope Imaging Spectrograph (CCD) was used to conduct bias and dark monitoring. Also, observations were made to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. As first mentioned in yesterday's report, two FHST updates failed immediately before the first iteration of this proposal began. There was no reported effect on these observations. Otherwise, the other observations were completed as planned, and no anomalies were reported. 1.4 Completed Three Sets STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.5 Completed WF/PC-2 6437 (Sins: Supernova Intensive Study - Cycle 6) The WF/PC-2 was used to make observations of the supernova LMC-SN1987 and its circumstellar matter which is located in the Large Magellanic Cloud. The observations were completed as planned, and no problems were noted. 1.6 Completed Six Sets of NIC/1/2/3 7726 (NICMOS One Orbit Parallel) The Near Infrared Camera (Cameras One, Two and Three) was used to make a one orbit parallel calibration observations. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Five Sets of STIS/CCD 7675 (STIS Non-Scripted Parallel Archive Proposal) The Space Telescope Imaging Spectrograph (CCD) was used in the parallel mode to obtain slitless spectroscopy with the G750L grating to search for emission line galaxies at intermediate to high redshift. The observations were completed as planned, and no anomalies were reported. 1.8 Completed NIC/2 7330 (The Fueling of Active Nucleii: A NICMOS Snapshot Survey) The Near Infrared Camera (Camera Two) was used to probe the dark matter distribution of NGC-157, far from the center, by observing intervening absorption from the MgII doublet (lambda 2800) in the spectrum of one background QSO located in a geometrically optimal position. HST observations are required because this feature is in a wavelength range not accessible from the ground, and the QSO magnitudes and expected strength of absorption make IUE observations unfeasible. As first mentioned in yesterday's report, two FHST updates failed immediately before this proposal began. There was no reported effect on the observations. Otherwise, there were no further anomalies. 1.9 Completed WF/PC-2 7476 (Sakurai's Nova-Like Object: Real-Time Monitoring of a Stellar Thermal Pulse) The WF/PC-2 was used to make images of Sakurai's nova-like object in order to do real-time monitoring of a stellar thermal pulse. The eruption of Sakurai's nova-like variable in Sagittarius presents a "once-in-a-lifetime" opportunity for real-time observations of a star undergoing a final helium thermal pulse. The observations were completed as planned, and no anomalies were reported. 1.10 Completed Two Sets of WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking ultraviolet observations running concurrent with prime science observations. The intent is to search for extra-galactic sources with flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. The observations were completed as planned, and no problems were reported. 1.11 Completed WF/PC-2/NIC/2 7450 (The Formation of Bulges of Spiral Galaxies: Central Visual-Infrared Color Gradients) The WF/PC-2 and the NIC (Camera Two) were both used to make observations of the bulge in spiral galaxy NGC-5838 in order to obtain high-resolution optical and near infrared images for a complete survey sample of bulges of spiral galaxies. The observations were completed as scheduled, and no anomalies were reported. 1.12 Completed Two Sets of STIS/CCD 7727 (STIS Non-Scripted Parallel Archive Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to obtain parallel slitless spectroscopy with the G750L grating to search for emission line galaxies at intermediate to high redshift. Field galaxy surveys have shown that at high redshift most galaxies have strong emission lines, and are compact and blue. The observations were completed as planned, and no anomalies were reported. 1.13 Completed NIC/2 7327 (Elliptical Galaxy Evolution at High Redshift: IR Imaging of 60 Distant Clusters) The Near Infrared Camera (Camera Two) was used to make infrared observations of high redshifted elliptical galaxies in galactic cluster CL1341+000. By measuring the structural properties of high redshift cluster elliptical galaxies, it is hoped to resolve the controversy regarding the age and formation history of bright cluster ellipticals. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed FGS 6880 (Parallax of RW Tri) The FGSs were used to make astrometric observations of the variable star RW-Tri. RW-Tri is an old nova for which a precise parallax would increase the validity of any astrophysical investigations. The observations were completed as planned, and no problems were reported. 1.15 Completed FOC/96 6667 (The Far-UV Evolution of Elliptical Galaxies) The Faint Object Camera (f/96) was used to make far-ultraviolet observations of the large elliptical galaxy G158-100. The far-UV spectra of nearby giant elliptical galaxies are characterized by a "UV rising branch", which varies greatly in strength from galaxy to galaxy. The observations were executes as planned, and no problems were reported. 1.16 Completed Three Sets of NIC/1/2/3 7729 (NICMOS Short One Orbit Parallel) The Near Infrared Camera (Cameras One, Two and Three) was used to obtain "JHK" imaging of sky regions while another instrument was prime. These observations were part of the short NICMOS one orbit parallel program. The observations were executed as scheduled, and no anomalies were noted. 1.17 Completed WF/PC-2 6836 (Light Curve of an Eclipsing Millisecond Pulsar) The WF/PC-2 was used to make observations of the eclipsing millisecond pulsar PSRJ2051-0 in order to obtain its light curve. This pulsar is an exciting new millisecond pulsar which is in a 2.38 hour circular orbit about a 0.03 solar mass companion which eclipses the pulsed radiation from the pulsar for 20 minutes each orbit. The companion is only the second eclipsing millisecond pulsar companion identified optically. The observations were executed as scheduled, and no anomalies were noted. 1.18 Completed NIC/2 7331 (Near-IR Properties of the Bulges of Spiral Galaxies) The Near Infrared Camera (Camera Two) was used to make infrared observations of the bulge in spiral galaxy NGC-4594. Near-IR data at HST resolution are essential for correcting for the effect of dust and will enable us to derive robust nuclear properties for the sample. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 12 Successful : 12 Scheduled Reacquisitions : 4 Successful : 4 2.2 FHST Updates: Scheduled : 32 Successful : 32 2.3 Operations Notes: The K10 bypass relay was closed at 191/1411Z. The SSR Interrupt Log was dumped at 192/0022Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science and SMOV operations. /DMH